# Learning Quadruped Locomotion Using Differentiable Simulation

### v2

@@ Abstract
While most recent advancements in legged robot control have been driven by model-free reinforcement learning, we explore the potential of differentiable simulation. Differentiable simulation promises faster convergence and more stable training by computing low-variant first-order gradients using the robot model, but so far, its use for legged robot control has remained limited to simulation. The main challenge with differentiable simulation lies in the complex optimization landscape of robotic tasks due to discontinuities in contact-rich environments, e.g., quadruped locomotion. This work proposes a new, differentiable simulation framework to overcome these challenges. The key idea involves decoupling the complex whole-body simulation, which may exhibit discontinuities due to contact, into two separate continuous domains. Subsequently, we align the robot state resulting from the simplified model with a more precise, non-differentiable simulator to maintain sufficient simulation accuracy. Our framework enables learning quadruped walking in minutes using a single simulated robot without any parallelization. When augmented with GPU parallelization, our approach allows the quadruped robot to master diverse locomotion skills, including trot, pace, bound, and gallop, on challenging terrains in minutes. Additionally, our policy achieves robust locomotion performance in the real world zero-shot. To the best of our knowledge, this work represents the first demonstration of using differentiable simulation for controlling a real quadruped robot. This work provides several important insights into using differentiable simulations for legged locomotion in the real world.
I Introduction
Many grazing animals, such as giraffes, lambs, and calves, can learn to stand and walk within minutes to hours after birth. On the other hand, most machine learning algorithms, particularly traditional model-free reinforcement learning (RL), need massive parallelization to achieve a stable walking policy. How can legged robots master walking in minutes with a few trials and errors?
Recent progress in legged robot control has largely been driven by GPU-accelerated massive simulation and reinforcement learning [15, 20, 24, 30, 22, 9, 39, 19]. Central to these advancements is the policy gradient algorithm [35, 34, 6]. The policy gradient can generally be expressed as the following
$\nabla_{\theta}J(\theta)\approx\mathbb{E}[\nabla_{\theta}\log\pi_{\theta}(u_{t}|x_{t})R_{t}(\bm{\tau})],$
where it stands as a zeroth-order approximation of the true gradient based on sampled trajectories $\bm{\tau}$. Here, $R_{t}(\bm{\tau})$ represents the sum of future rewards. Despite impressive real-world performance, this approach to gradient estimation is known for its high variance. Consequently, several additional strategies are required for stable training, such as using a clipped surrogate objective [31], increasing the number of samples, designing task-specific learning curriculums, employing distributed initialization techniques, and engineering suitable reward functions.
As a result, RL struggles with complex tasks where data generalization is slow, like vision-based navigation. Despite the need for more effective learning algorithms, progress has been limited. Instead, research leans towards imitation learning, utilizing existing datasets or expert demonstrations for more consistent learning outcomes. Notably, the learning-by-cheating approach [4] is a popular imitation learning framework: first, you train an RL teacher with privileged information and then, you derive a student policy based on visual inputs. However, the success of this approach heavily depends on the quality of expert demos and lacks generalizability and adaptability to new tasks.
In robotics, leveraging well-established knowledge about robot dynamics can enable the construction of first-order gradient estimations, which typically exhibit significantly lower variance than their zeroth-order counterparts and hold great potential for more stable training and faster convergence. Recently, policy training using first-order gradient has been notably advanced through differentiable simulation [8, 14, 36, 29, 37]; these works have shown promising results in reducing both the number of simulation samples and total training time compared to zeroth-order methods.
@@ I Introduction
Figure 1: Graphical model for policy learning using differentiable simulation, in which both the loss function and the dynamics are differentiable.
However, the optimization landscape in robotic tasks is inherently complex, characterized by nonlinearities, non-smooth and discontinuous dynamics, non-convexities, and long temporal horizons. Additionally, deploying real-world robots introduces additional challenges, such as the sim-to-real gap, noisy state estimation, and system delay. Consequently, policy training through differentiable simulation can suffer from issues like local minimum and gradient vanishing/explosion. While differentiable simulation offers a promising avenue, its effectiveness in significantly accelerating policy learning in contact-rich environments and its applicability to real-world scenarios like dynamic quadruped control remains an open challenge in robotics.
This work set out to explore the potential of differentiable simulation in accelerating policy training within contact-rich scenarios like dynamic legged locomotion. Our goal is to establish an effective policy training framework that leverages the robot model for precise gradient computation while mitigating the difficulties posed by complex, discontinuous optimization landscapes. Moreover, we intend to evaluate the strengths and weaknesses of this approach in comparison to zeroth-order methods, particularly model-free policy gradients. We aim to provide useful insights for developing more efficient and robust methods that benefit from both worlds.
@@
Contribution: We show that differentiable simulation offers considerable advantages over model-free reinforcement learning for policy training in legged locomotion. Notably, we demonstrate that a single robot, without the need for parallelization, can quickly learn to walk within minutes using our approach. Leveraging the advantage of GPUs for parallelized simulation, our robot learns diverse walking skills over challenging terrain in minutes. Specifically, we train a quadruped robot to walk with different gait patterns, including trot, pace, bound, and gallop, and with varying gait frequencies. While model-free RL can also achieve similar performance given sufficient parallelization, our approach requires much less data and yields better performance.
More importantly, we show that the policy trained via differentiable simulation can be transferred to the real world directly without fine-tuning. To the best of our knowledge, this work represents the first demonstration of using differentiable simulation for controlling a real quadruped robot. This highlights the effectiveness of differentiable simulation for real-world applications in dynamic legged locomotion scenarios.
@@
The key to our approach is a novel policy training framework that combines the smooth gradients obtained from a simplified dynamics model for efficient backpropagation with the high fidelity of a more complex, non-differentiable simulator for accurate forward simulation. 1) Decoupling Simulation Spaces: Rather than handling the whole body dynamics as a unified differentiable simulator, which may encounter non-smoothness due to contact interactions. We proposed to separate the simulation into two distinct spaces: the floating base space and the joint space. For the robot body, we employ an approximation using single rigid-body dynamics, which offers a continuous and effective representation of the floating base (the robot body). 2) PD Control as A Differentiable Layer: The simulation within the joint space benefits from the inherent differentiability and smoothness of a Proportional-Derivative (PD) controller. This allows us to treat the PD control as an explicit, differentiable layer within our simulation framework. 3) Alignment with Non-Differentiable Simulators: To address potential inaccuracies arising from our simplified rigid-body dynamics model, we incorporate a more precise, non-differentiable simulator. This non-differentiable simulator can simulate complex contact dynamics and is used to align our simplified model, thereby ensuring that our training pipeline remains grounded in realistic dynamics.
II Related Work
II-A Model Predictive Control
Model predictive control (MPC) has a long history in legged locomotion [18, 16]. MPC relies on fast online optimization and an accurate dynamics model to control the robot. To trade-off between model accuracy and computation speed, many methods use a reduced optimization model, such as single rigid-body dynamics [7, 11, 2]. Research has also studied control using more complex formulations, including centroidal [10] and whole-body dynamics [26]. Despite their success, MPC lacks the flexibility in its optimization objective design and robustness when facing disturbances or unexpected scenarios. Empirical evidence suggests that MPC tends to underperform when compared to data-driven approaches such as deep reinforcement learning.
II-B Reinforcement Learning
Deep reinforcement learning has emerged as a dominant approach for developing control policies in legged locomotion [15, 20, 24, 19]. Recently, an important advancement in this area has been the introduction of IsaacGym [21], which is a GPU-accelerated simulator. IsaacGym can dramatically reduce the time required for dynamics simulation and policy training, enabling robots to learn to walk on flat ground in minutes [30]. This advancement has significantly accelerated the pace of research in legged locomotion, opening new avenues to address increasingly complex challenges [38, 1, 9, 22, 39].
Nevertheless, it is important to highlight that the progress in applying reinforcement learning to legged locomotion is primarily driven by the enhanced computational capabilities provided by modern GPUs rather than substantial breakthroughs in the underlying algorithms. Consequently, in scenarios where data collection cannot be accelerated through computational means, researchers may resort to alternative strategies such as “learning-by-cheating” to circumvent these limitations. This highlights a gap in the current approach to deep RL in legged locomotion, suggesting a need for innovative algorithmic developments that do not solely rely on computational advancements.
II-C Differentiable Simulation
Differentiable simulation has recently gained momentum, thanks to the development of differentiable physics simulators [8, 5, 12] and flexible automatic differentiation tools like Jax, DiffTaichi, Pytorch [3, 13, 27]. In principle, policy training through differentiable simulation allows better convergence by replacing the zeroth-order gradient estimation of a stochastic objective with a more accurate estimate based on first-order gradients [33]. However, challenges such as the noisy optimization landscape and issues of exploding or vanishing gradients in long-horizon tasks render first-order gradient methods less effective.
Several approaches have been developed to tackle these challenges. For instance, SHAC [36] addresses the gradient issue by truncating trajectories into several smaller segments, which helps manage exploding or vanishing gradients. On the other hand, PODS [25] takes advantage of second-order gradients. This approach facilitates monotonic policy improvement and promises faster convergence than first-order methods. However, PODS’ effectiveness relies on very accurate computation of second-order derivatives from the differentiable simulator, making it sensitive to the precision of these derivatives. Another notable example is DiffMimic [29], which focuses on mimicking motion trajectories for physically simulated characters.
Research in differentiable simulation has mainly been focused on speeding up policy optimization using first-order gradients, alleviating local minima problems, addressing gradient vanishing/explosion issues, or developing control policies for simulated characters. They show that differentiable simulation is a promising direction for improving sample efficiency. However, the capability of differentiable simulation to substantially speed up policy learning in contact-rich scenarios with physical interactions, as well as their practical applicability to real-world situations, such as quadruped locomotion, continues to be a significant challenge in robotics.
Our research distinguishes itself from earlier studies by presenting a novel perspective that leads to an effective solution tailored for real-world robotic applications.
III Methodology
Figure 2: System overview of legged robot control using differentiable simulation. Our approach leverages the differentiability and smoothness of a single rigid-body dynamics for dynamics simulation. Additionally, we treat PD control as a differential layer in our computation graph. Finally, we use the state from a more accurate, non-differentiable simulator (IsaacGym) to align the simulated state in the simple rigid-body simulation.
III-A Overview
Fig. 2 shows our training setup that employs differentiable simulation. The fundamental idea is that the main body of a legged robot can be modeled using differentiable and smooth single rigid-body dynamics, while the simulation of the robot’s four legs can be approached separately from its main body. We use a Proportional-Derivative (PD) controller to connect the legs dynamics with the main body. In the forward pass, this PD controller converts joint angles into torques, which then produce the contact forces needed for the robot’s rigid-body simulation. During backpropagation, the PD controller functions as a differentiable component, enabling gradients to be propagated back through the system. Finally, we use non-differentiable whole-body dynamics simulation to align the state produced by our simplified rigid-body dynamics.
III-B Problem Formulation
@@
We formulate legged robot control as an optimization problem. The robot is modeled as a discrete-time dynamical system, characterized by continuous state and control input spaces, denoted as $\mathcal{X}$ and $\mathcal{U}$, respectively. At each time step $k$, the system state is $x_{k}\in\mathcal{X}$, and the corresponding control input is $u_{k}\in\mathcal{U}$. An observation $o_{k}\in\mathcal{O}$ is generated at each time step based on the current state $x_{k}$ through a sensor model $h:\mathcal{X}\rightarrow\mathcal{O}$, such that $o_{k}=h(x_{k})$. The system’s dynamics are governed by the function $f:\mathcal{X}\times\mathcal{U}\rightarrow\mathcal{X}$, which describes the time-discretized evolution of the system as $x_{k+1}=x_{k}+\Delta t\cdot f(x_{k},u_{k})$. The discrete-time instants are $t_{k}=\Delta t\cdot k$, where $k$ ranges from 0 to $N$, thereby establishing a finite time horizon for the control problem. At each time step $k$, the robot receives a cost signal $l_{k}=l(x_{k},u_{k})$, which is a function of the current state $x_{k}$ and the control input $u_{k}$.
The control policy is represented as a deterministic, differentiable function such as a neural network $u_{k}=\pi_{\theta}(o_{k})$. The neural network takes the observation $o_{k}$ as input and outputs the control input $u_{k}$. The optimization objective is to find the optimal policy parameters $\theta^{\ast}$ by minimizing the total loss via gradient descent:
@@
$\theta$ $\leftarrow\theta-\alpha\nabla_{\theta}\mathcal{L}_{\theta}$ (2)
@@
III-C Forward Simulation
Legged robots are characterized by complex dynamics resulting from their multi-body configuration and the contact between their body and the environment. The contacts introduce discontinuities in the dynamics, making first-order gradient-based optimization algorithms difficult. An important aspect of our approach involves separating the simulation of the robot’s body from that of its joints. This design choice allows a simple implementation of the differentiable simulator.
First, we represent the robot’s main body using single rigid-body dynamics. Single rigid-body dynamics have been shown to be useful for dynamic locomotion using online gradient-based optimization [7, 11]. The single rigid-body model offers a continuous representation of the robot base dynamics, avoiding the complex optimization landscape introduced by contacts. We develop our differentiable simulation using the single rigid-body dynamics, which is expressed as follows,
@@ III-C Forward Simulation
$\dot{\mathbf{q}}_{WB}$ $=\frac{1}{2}\Lambda(\bm{\omega}_{B})\cdot\mathbf{q}_{WB}$ $\dot{\bm{\omega}}_{B}$ $=\mathbf{I}^{-1}\left(\bm{\eta}-\bm{\omega}_{B}\times(\mathbf{I}\bm{\omega}_{B})\right)$
In this approximation, the state of our system is
$\mathbf{x}=[\mathbf{p},\mathbf{q},\mathbf{v},\bm{\omega}],$
where $\mathbf{p}_{\text{WB}}\in\mathbb{R}^{3}$ is the position and $\mathbf{v}_{\text{WB}}\in\mathbb{R}^{3}$ is the linear velocity of the center of mass in the world frame ${W}$, We use a unit quaternion $\mathbf{q}_{\text{WB}}$ to represent the orientation of the body in the world frame and use $\bm{\omega}_{\text{B}}$ to denote the body rates in the body frame ${B}$. Here, $\mathbf{\Lambda}(\bm{\omega}_{B})$ is a skew-symmetric matrix. The control inputs are the ground reaction force $\mathbf{f}_{i}$ from the legs.
To simulate the rigid-body dynamics, it is important to know the ground reaction force. One option for the control policy design is to output the ground reaction force directly, similar to the MPC design [7, 11]. Another option is to control the robot in the joint space, e.g., output desired joint position, which allows more control authority for the policy and adaptive behavior. In this case, it is required to convert the neural network output (joint position) into the control input of the single rigid-body model, which is the ground reaction force. This conversion is achieved using a PD controller for forward propagation,
$\bm{\tau}=\mathbf{k_{p}(q^{\text{ref}}-q)}+\mathbf{k_{d}(\dot{q}^{\text{ref}}-\dot{q})}$ (3)
which calculates the required motor torque $\bm{\tau}$. Here, $\mathbf{k_{p}}$ and $\mathbf{k_{d}}$ are fixed gains, $\mathbf{q^{\text{ref}}}$ and $\mathbf{q}$ are the reference joint position and the current joint position separately, and $\mathbf{\dot{q}^{\text{ref}}}$ and $\mathbf{\dot{q}}$ are the reference joint velocity and the current joint velocity respectively.
Subsequently, the motor torques are then converted to ground reaction forces $\mathbf{f}$ using the foot Jacobian $\mathbf{J}$:
$\mathbf{f}=(\mathbf{J}^{T})^{-1}\bm{\tau}.$
The continuous nature of PD control enables the backpropagation of policy gradients. Consequently, our simulation framework treats the PD controller as an explicit, differentiable layer.
III-D Backpropagation Through Time

### v3

@@ Abstract
This work explores the potential of using differentiable simulation for learning quadruped locomotion. Differentiable simulation promises fast convergence and stable training by computing low-variance first-order gradients using robot dynamics. However, its usage for legged robots is still limited to simulation. The main challenge lies in the complex optimization landscape of robotic tasks due to discontinuous dynamics. This work proposes a new differentiable simulation framework to overcome these challenges. Our approach combines a high-fidelity, non-differentiable simulator for forward dynamics with a simplified surrogate model for gradient back-propagation. This approach maintains simulation accuracy by aligning the robot states from the surrogate model with those of the precise, non-differentiable simulator. Our framework enables learning quadruped walking in simulation in minutes without parallelization. When augmented with GPU parallelization, our approach allows the quadruped robot to master diverse locomotion skills on challenging terrains in minutes. We demonstrate that differentiable simulation outperforms a reinforcement learning algorithm (PPO) by achieving significantly better sample efficiency while maintaining its effectiveness in handling large-scale environments. Our method represents one of the first successful applications of differentiable simulation to real-world quadruped locomotion, offering a compelling alternative to traditional RL methods.
Video: https://youtu.be/weNq_w715xM
Keywords: Differentiable Simulation, Legged Locomotion
1 Introduction
@@
Figure 1: Graphical model for policy learning using differentiable simulation.
Traditional model-free reinforcement learning (RL) often requires extensive parallelization to achieve a stable walking policy for legged robots [1]. How can legged robots master walking quickly with a few trials and errors? Recent progress in legged robot control has been largely driven by combining model-free RL with massively parallelized simulation [1, 2, 3, 4, 5, 6, 7]. Central to these advancements is the policy gradient algorithm [8, 9], where the gradient of a control policy can be calculated via $\nabla_{\theta}J(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta}}[R(\tau)\nabla_{\theta}\log p(\tau;\theta)]$; which is a zeroth-order estimate of the true gradient of $J$ based on sampled trajectories following the policy $\pi_{\theta}$. Here, $R(\tau)$ is the reward of a trajectory $\tau$. Such zeroth-order gradient estimate provides a powerful and general framework for robot control since it can handle non-differentiable optimization objectives and discontinuous dynamics. However, the gradient has high variances. Consequently, several additional strategies are required for stable training, such as using a clipped surrogate objective [10] and increasing the sample size.
In robotics, leveraging well-established knowledges about robot dynamics can enable the construction of first-order gradient estimates $\nabla_{\theta}J(\theta)=\nabla_{\theta}R(\tau)$, which typically exhibit significantly lower variance than their zeroth-order counterparts and hold great potential for more stable training and faster convergence. Recently, policy training using first-order gradient has been notably advanced through differentiable simulation [11, 12, 13, 14, 15]; these works have shown promising results in reducing both the number of simulation samples and total training time compared to zeroth-order methods. Differentiable simulation combines the advantages of model-based control and data-driven learning. Similar to numerical optimal control methods, such as shooting techniques, differentiable simulation utilizes robot dynamics to simulate trajectories and iteratively optimize decision variables using analytical gradients derived from the robot model. Like reinforcement learning, differentiable simulation benefits from large-scale simulations and deep learning to train neural network control policies that map observations directly to control actions. Despite its advantages, differentiable simulation faces several challenges, including discontinuous dynamics, complex optimization landscapes, and issues like exploding or vanishing gradients in long-horizon tasks, which can limit the effectiveness of first-order gradient methods.
@@
This work investigates the potential of training policies through surrogate models within differentiable simulation, specifically focusing on quadruped locomotion. We show that differentiable simulation offers considerable advantages over model-free RL for policy training in legged locomotion. Notably, we demonstrate that a single robot, without the need for parallelization can quickly learn to walk within minutes in simulation using our approach. Leveraging the advantage of GPUs for parallelized simulation, our robot learns diverse walking skills over challenging terrain in minutes. Specifically, we train a quadruped robot to walk with different gait patterns, including trot, pace, bound, and gallop, and with varying gait frequencies. While model-free RL can also achieve similar performance given sufficient parallelization, our approach requires much less data, achieving significantly better sample efficiency while maintaining its effectiveness in handling large-scale environments. More importantly, we show that the policy trained via differentiable simulation can be transferred to the real world directly without fine-tuning. This work presents one of the earliest demonstrations of using differentiable simulation to train control policies for real quadruped robots, highlighting the potential of differentiable simulation for real-world applications.
@@
Contribution: The key to our approach is a novel policy training framework that combines the smooth gradients obtained from a simple surrogate dynamics model for efficient backpropagation with the high fidelity of a more complex, non-differentiable simulator for accurate forward simulation. Instead of simulating the quadruped robot using whole-body dynamics, which by definition has discontinuities due to contact, we propose separating the simulation into its floating base space and its joint space. For the robot body, we employ an approximation using single rigid-body dynamics, which offers a continuous and effective representation of the robot body. To address simulation errors arising from our simplified rigid-body dynamics model, we incorporate a more precise, non-differentiable simulator. This non-differentiable simulator can simulate complex contact dynamics and is used to align the state in our simplified model, thereby ensuring that our training pipeline remains grounded in realistic dynamics. Figure 2 provides an overview of our system.
2 Methodology
Figure 2: System overview of learning quadruped locomotion using differentiable simulation. Our approach decouples the robot dynamics into two separate spaces: joint and floating base spaces. We leverage the differentiability and smoothness of a single rigid-body dynamics for the robot’s main body, which takes the ground reaction force from its legs as the control inputs. We use the state from a non-differentiable simulator (IsaacGym) to align the state in the differentiable simulation.
2.1 Problem Formulation
@@
We formulate legged robot control as an optimization problem. The robot is modeled as a discrete-time dynamical system, characterized by continuous state and control input spaces, denoted as $\mathcal{X}$ and $\mathcal{U}$, respectively. At each time step $k$, the system state is $x_{k}\in\mathcal{X}$, and the corresponding control input is $u_{k}\in\mathcal{U}$. An observation $o_{k}\in\mathcal{O}$ is generated at each time step based on the current state $x_{k}$ through a sensor model $h:\mathcal{X}\rightarrow\mathcal{O}$, such that $o_{k}=h(x_{k})$. The system’s dynamics are governed by the function $f:\mathcal{X}\times\mathcal{U}\rightarrow\mathcal{X}$, which describes the time-discretized evolution of the system as $x_{k+1}=f(x_{k},u_{k})$. At each time step $k$, the robot receives a cost signal $l_{k}=l(x_{k},u_{k})$, which is a function of the current state $x_{k}$ and the control input $u_{k}$. The control policy is represented as a deterministic, differentiable function such as a neural network $u_{k}=\pi_{\theta}(o_{k})$. The neural network takes the observation $o_{k}$ as input and outputs the control input $u_{k}$. The optimization objective is to find the optimal policy parameters $\theta^{\ast}$ by minimizing the total loss via gradient descent
@@
$\theta\leftarrow\theta-\alpha\nabla_{\theta}\mathcal{L}_{\theta},$ (2)
@@
2.2 Forward Simulation
We represent the robot’s main body using single rigid-body dynamics. Single rigid-body dynamics have been shown to be useful for dynamic locomotion using model predictive control [16, 17]. The single rigid-body model offers a continuous representation of the robot base dynamics, avoiding the complex optimization landscape introduced by contacts. We develop our differentiable simulation using the single rigid-body dynamics, which is expressed as follows,
@@ 2.2 Forward Simulation
$\dot{\mathbf{q}}_{WB}$ $=\frac{1}{2}\Lambda(\bm{\omega}_{B})\cdot\mathbf{q}_{WB}$ $\dot{\bm{\omega}}_{B}$ $=\mathbf{I}^{-1}\left(\bm{\eta}-\bm{\omega}_{B}\times(\mathbf{I}\bm{\omega}_{B})\right).$
In this approximation, the state of our system is $\mathbf{x}=[\mathbf{p},\mathbf{q},\mathbf{v},\bm{\omega}],$ where $\mathbf{p}_{\text{WB}}\in\mathbb{R}^{3}$ is the position and $\mathbf{v}_{\text{WB}}\in\mathbb{R}^{3}$ is the linear velocity of the center of mass in the world frame ${W}$, We use a unit quaternion $\mathbf{q}_{\text{WB}}$ to represent the orientation of the body in the world frame and use $\bm{\omega}_{\text{B}}$ to denote the body rates in the body frame ${B}$. Here, $\mathbf{I}$ is the robot’s inertia tensor, $\bm{\eta}$ is the body torque, and $\mathbf{\Lambda}(\bm{\omega}_{B})$ is a skew-symmetric matrix. The control inputs are the ground reaction force $\mathbf{f}_{i}$ from the legs that have contacts.
The ground reaction forces are required to simulate the rigid-body dynamics. One option for the control policy design is to output the ground reaction force directly, similar to the MPC design [16, 17]. Another option is to control the robot in the joint space, e.g., output desired joint position, which allows more control authority for the policy and adaptive behavior. Our policy maps observations to the desired joint position $\mathbf{q}^{\text{ref}}$, while assuming zero joint velocity, i.e., $\mathbf{\dot{q}}^{\text{ref}}=\mathbf{0}$. In this case, the neural network output (joint position) is required to be converted into the control input of the single rigid-body model, which is the ground reaction force. This conversion is achieved using a PD controller for forward propagation
$\bm{\tau}=\mathbf{k_{p}(q^{\text{ref}}-q)}+\mathbf{k_{d}(\dot{q}^{\text{ref}}-\dot{q})},$ (3)
which calculates the required motor torque $\bm{\tau}$. Here, $\mathbf{k_{p}}$ and $\mathbf{k_{d}}$ are fixed gains, $\mathbf{q^{\text{ref}}}$ and $\mathbf{q}$ are the reference joint position and the current joint position separately, and $\mathbf{\dot{q}^{\text{ref}}}$ and $\mathbf{\dot{q}}$ are the reference joint velocity and the current joint velocity respectively. Subsequently, the motor torques are then converted to ground reaction forces $\mathbf{f}$ using the foot Jacobian $\mathbf{J}$: $\mathbf{f}=(\mathbf{J}^{T})^{-1}\bm{\tau}.$ The continuous nature of PD control enables the backpropagation of policy gradients. Consequently, our simulation framework treats the PD controller as a differentiable layer.
2.3 Backpropagation Through Time